# Multi-LoRA Composition for Image Generation

### v1

@@ 3.2 Results on ComposLoRA
3.3 Analysis
To enhance our understanding of the proposed methods, we further investigate the following critical questions:
@@ 3.3.1 Do Specific Image Styles Favor Different Methods?
To explore this, we separately evaluate the performance of methods on realistic and anime-style subsets within ComposLoRA. The win rate results, presented in Figure 3, reveal distinct tendencies for each method. Our observations reveal that, while LoRA-s may not excel in image quality compared to LoRA-c, it demonstrates comparable performance in this dimension within the realistic style subset, while maintaining a significant edge in composition quality. In contrast, in the anime-style subset, LoRA-c, shows a performance on par with LoRA-s in composition quality, while notably surpassing it in image quality. These findings suggest that LoRA-S is more adept at composing elements in realistic-style images, whereas LoRA-C shows a stronger performance in anime-style imagery.
@@ 3.3.2 How Does the Step Size and Order of LoRA Activation Affect LoRA Switch?
To identify the optimal configuration for LoRA Switch, we examine the influence of two crucial hyperparameters: the sequence in which LoRAs are activated and the interval between each activation. Our findings, depicted in Figure LABEL:fig:switch_step, show that overly frequent switching, such as changing LoRAs at every denoising step, leads to distortions in generated images and suboptimal performance. The efficiency of the LoRA Switch improves progressively with increased step size, reaching peak performance at $\tau$ = 5. Consequently, we adopt this step size in our experiments.
Moreover, our analysis underscores that the initial choice of LoRA in the activation sequence clearly influences overall performance, while alterations in the subsequent order have minimal impact. Activating the character LoRA first leads to the best performance, as demonstrated in Figure LABEL:fig:switch_order. In contrast, starting with clothing, background, or object LoRAs yields results comparable to a completely randomized sequence. Notably, beginning with the style LoRA leads to a noticeable performance drop, even falling slightly below a random order. This observation underlines the critical role of prioritizing core image elements in the initial stage of the generation process to enhance both the image and compositional quality for LoRA Switch.
Figure 4: Positional bias analysis for comparative evaluation uses GPT-4V. In each subfigure, the left side of the orange line compares LoRA-s with Merge, and the right side contrasts LoRA-c with Merge. “Merge First” indicates that the image produced by LoRA Merge is the first image input to GPT-4V during the comparative evaluation.
@@ 3.3.3 Does GPT-4V Exhibit Bias as an Evaluator?
While GPT-4V has demonstrated utility in evaluating various image generation tasks (Lin et al., 2023; Zhang et al., 2023b), our analysis uncovers a notable positional bias in its comparative evaluations. We investigate this potential bias by swapping the positions of images generated by different methods before inputting them to GPT-4V, and the results are illustrated in Figure 4. In the comparison of LoRA-s versus LoRA Merge, when the image generated by Merge is presented first (“Merge First”), the win rate for LoRA-s in composition quality stands at 60%. However, this win rate declines to 51% when LoRA-s’s image is the first input (“LoRA-S First”). Similarly, LoRA-c’s win rate decreases from 52% to 42%, suggesting that GPT-4V tends to favor the first image input in terms of composition quality. Intriguingly, the opposite trend is observed in image quality, where the second image tends to receive a higher score. These results indicate a significant positional bias in GPT-4V’s evaluation, varying with the dimension and the position of the images. To mitigate this bias in our study, the comparative evaluation results reported in this paper are averaged across both input orders.
4 Related Work
4.1 Composable Text-to-Image Generation
Composable image generation, a key aspect of digital content customization, involves creating images that adhere to a set of pre-defined specifications (Liu et al., 2023). Existing research in this domain primarily focuses on the following approaches: enhancing compositionality with scene graphs or layouts (Johnson et al., 2018; Yang et al., 2022; Gafni et al., 2022), modifying the generative process of diffusion models to align with the underlying specifications (Feng et al., 2023; Huang et al., 2023c, b), or composing a series of independent models that enforce desired constraints (Du et al., 2020; Liu et al., 2021; Nie et al., 2021; Liu et al., 2022; Li et al., 2023; Du et al., 2023).
However, these methods typically operate at the concept level, where generative models excel in creating images based on broader categories or general concepts. For example, a model might be prompted to generate an image of “a woman wearing a dress”, and can adeptly accommodate variations in the textual description, such as changing the color of the dress. Yet, they struggle to accurately render specific, user-defined elements, like lesser-known characters or unique dress styles. Another line of work that can compose user-defined objects into images includes Huang et al. (2023c); Ruiz et al. (2023). However, these methods require extensive fine-tuning and do not perform well on multiple objects. Therefore, we introduce learning-free instance-level composition approaches utilizing LoRA in this paper, enabling the precise assembly of user-specified elements in image generation.
4.2 LoRA-based Manipulations
5 Conclusion
Impact Statements
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Appendix
@@
Table 3: Detailed descriptions of each LoRA in the ComposLoRA.
LoRA Category Trigger Words Source
Anime Style Subset
Zero Suit (Metroid) Clothing zero suit, blue gloves, high heels Link
Hand-drawn Style Style lineart, hand-drawn style Link
Auroral Background Background auroral, starry sky, outdoors Link
Toast Object toast, toast in mouth Link
Realistic Style Subset
Japanese Film Color Style Style film overlay, film grain Link
Bright Style Style bright lighting Link
Forest Background Background slg, river, forest Link
Umbrella Object transparent umbrella Link
Bubble Gum Object blow bubble gum Link

### v2

@@ 4.3.1 Do Specific Image Styles Favor Different Methods?
To explore the impact of image style, we separately evaluate the performance of methods on realistic and anime-style subsets within ComposLoRA. The win rate results, presented in Figure 4, reveal distinct tendencies for each method.
Our observations reveal that, while LoRA-s may not excel in image quality compared to LoRA-c, it demonstrates comparable performance in this dimension within the realistic style subset, while maintaining a significant edge in composition quality. In contrast, in the anime-style subset, LoRA-c, shows a performance on par with LoRA-s in composition quality, while notably surpassing it in image quality. These findings suggest that LoRA-S is more adept at composing elements in realistic-style images, whereas LoRA-C shows a stronger performance in anime-style imagery.
@@ 4.3.2 How Does the Step Size and Order of LoRA Activation Affect LoRA Switch?
To identify the optimal configuration for LoRA Switch, we examine the influence of two crucial hyperparameters: the sequence in which LoRAs are activated and the interval between each activation. Our findings, depicted in Figure 5(a), show that overly frequent switching, such as changing LoRAs at every denoising step, leads to distortions in generated images and suboptimal performance. The efficiency of the LoRA Switch improves progressively with increased step size, reaching peak performance at $\tau$ = 5.
(a) Steps to Switch
(b) Activation Order
Figure 5: Analysis of the number of denoising steps to switch LoRA and the activation order for LoRA Switch. In Figure 5(b), “Character” indicates that the character LoRA is activated first, with the rest being activated randomly.
Moreover, our analysis underscores that the initial choice of LoRA in the activation sequence clearly influences overall performance, while alterations in the subsequent order have minimal impact. Activating the character LoRA first leads to the best performance, as demonstrated in Figure 5(b). In contrast, starting with clothing, background, or object LoRAs yields results comparable to a completely randomized sequence. Notably, beginning with the style LoRA leads to a noticeable performance drop, even falling slightly below a random order. This observation underlines the critical role of prioritizing core image elements in the initial stage of the generation process to enhance both the image and compositional quality for LoRA Switch.
While the step size for switching LoRAs proves to be a crucial factor in achieving optimal performance in our experiments, we also explore the potential of dynamic strategies for step size adjustment throughout the denoising process. Specifically, we design and evaluate three strategies for dynamically adjusting the step size in LoRA-Switch:
  • Incremental Strategy: The step size gradually increases from $\tau=3$ to $\tau=5$ throughout the denoising process.
  • Decremental Strategy: The step size gradually decreases from $\tau=5$ to $\tau=3$ as the denoising process progresses.
  • Warm-up Strategy: During the initial 50% of the denoising process, the step size increases from $\tau=3$ to $\tau=5$ and remains constant at $\tau=5$ for the remaining denoising steps.
Table 3: Performance comparison of dynamic strategies for LoRA-S.
$\tau=3$ $\tau=4$ $\tau=5$ $\tau=6$ Incremental (3 $\rightarrow$ 5) Decremental (5 $\rightarrow$ 3) Warm-up
Composition Quality 55 57 59 58 57 54 58
Image Quality 29 33 34 33 33 31 34
Table 3 shows the results of these strategies. Neither the Incremental nor the Warm-up strategies significantly improve performance compared to using a fixed step size. The Decremental strategy, on the other hand, results in comparatively worse performance, highlighting that switching LoRAs too frequently in the latter stages of denoising is detrimental to image quality. The fixed step size of $\tau=5$ yields the best performance. Consequently, we adopt this fixed step size in our experiments.
@@ 4.3.3 Does GPT-4V Exhibit Bias as an Evaluator?
While GPT-4V has demonstrated utility in evaluating various image generation tasks (Lin et al., 2023; Zhang et al., 2023b), our analysis uncovers a notable positional bias in its comparative evaluations. We investigate this potential bias by swapping the positions of images generated by different methods before inputting them to GPT-4V, and the results are illustrated in Figure 6.
Figure 6: Positional bias analysis for GPT-4V-based evaluation. In each subfigure, the left side of the orange line compares LoRA-s with Merge, and the right side contrasts LoRA-c with Merge. “Merge First” indicates that the image produced by LoRA Merge is the first image input during the comparative evaluation.
In the comparison of LoRA-s versus LoRA Merge, when the image generated by Merge is presented first (“Merge First”), the win rate for LoRA-s in composition quality stands at 60%. However, this win rate declines to 51% when LoRA-s’s image is the first input (“LoRA-S First”). Similarly, LoRA-c’s win rate decreases from 52% to 42%, suggesting that GPT-4V tends to favor the second image input in terms of composition quality. Intriguingly, the opposite trend is observed in image quality, where the second image tends to receive a higher score. These results indicate a significant positional bias in GPT-4V’s evaluation, varying with the dimension and the position of the images. To mitigate this bias in our study, the comparative evaluation results reported in this paper are averaged across both input orders.
4.4 More Visual Examples
To demonstrate the effectiveness of our methods in composing varying numbers of LoRAs and under different image styles, we provide additional visual examples in Figures 7 – 10.
@@ 4.4 More Visual Examples
5 Further Discussions
5.1 Limitations
Based on our experiments, the primary limitation of the proposed methods is the efficiency issue with LoRA-c. This method can introduce $(k-1)\times$ additional computational cost, where $k$ is the number of merged LoRAs. This is due to LoRA-c merging each LoRA with the base model to calculate scores, which are then averaged. The inherent design of LoRA prevents the pre-computation of the base model. To address this, we propose two potential solutions: 1) Integrating advanced techniques with fewer denoising steps, and 2) a combination of LoRA-s and LoRA-c.
5.1.1 Integration of LCM and LCM-LoRA
Recently, several algorithms have been developed that require only 2-8 denoising steps to generate high-quality images. We select LCM (Luo et al., 2023a) and LCM-LoRA (Luo et al., 2023b) to test if our approaches can integrate smoothly with these algorithms.
Table 4: Results for integrating LCM and LCM-LoRA.
Integrated Technique Method Win (%) Tie (%) Lose (%) Inference Steps
None LoRA Switch 43 34 23 200
LoRA Composite 33 51 16 200
LCM LoRA Switch 45 36 19 4
LoRA Composite 39 43 18 4
LCM-LoRA LoRA Switch 42 39 19 8
LoRA Composite 44 41 15 8
First, we conduct experiments on 2 LoRAs using the same setup as described in the main text, with results shown in Table 4. Remarkably, our methods not only outperform the baseline but also show even greater advantages when integrated with these inference-accelerating techniques. This adaptation significantly reduces the required number of denoising steps to 4-8, effectively addressing the increased computational demands of LoRA-c. Consequently, the generation times across all three methods are now comparably short, taking only a few seconds.
To further explore the potential of integrating our method with fewer denoising steps, we conduct additional experiments using LCM-LoRA with an increased number of LoRAs. These experiments aim to provide deeper insights into how our approach performs as the complexity of multi-LoRA composition increases. The results, shown in Table 5, reflect the composition quality scores under this setup. For these experiments, we use 8 denoising steps for 2 and 4 LoRAs, and 9 inference steps for 3 LoRAs, switching to the next LoRA every 3 steps for LoRA-s.
Table 5: Composition quality scores for different methods when using LCM-LoRA with more LoRAs.
2 LoRAs 3 LoRAs 4 LoRAs
LoRA Merge 7.52 5.16 3.49
LoRA Switch 8.08 6.39 5.08
LoRA Composite 8.13 6.07 4.53
Our methods exhibit a more pronounced advantage over the baseline in this fewer-steps setting. The improvement is particularly noticeable when the number of LoRAs increases, demonstrating the robustness of our approach even when integrated with models requiring fewer denoising steps. However, it is important to note that all methods experience a substantial drop in absolute scores when combined with models like LCM-LoRA that employ fewer denoising steps. For instance, the composition quality score for LoRA-c with 5 LoRAs is initially 6.56 for 200 denoising steps, but with 4 LoRAs in this setting, the score drops to 5.08. This finding suggests that despite the improved performance of our method, multi-LoRA composition remains a challenging task, especially when fewer denoising steps are used, even with the latest integration techniques.
5.1.2 Combination of LoRA-s and LoRA-c
To further enhance efficiency, we propose combining LoRA-s and LoRA-c. LoRA-s activates at the LoRA stage before each denoising step, while LoRA-c is applied to the CFG during the denoising process. These design principles are complementary. A practical integration method involves selecting a subset of LoRAs to activate (ranging from one to all) for each denoising step, following the LoRA-s strategy. This subset would then utilize all its LoRAs during the denoising phase, adhering to the LoRA-c strategy. For LCM and other related applications, combining LoRA-s and LoRA-c can enhance efficiency without additional modifications. For example, in a 1-step scenario, all LoRAs can be activated and applied through CFG (as per LoRA-c). In a 2-step scenario, half of the LoRAs can be activated at each step and then applied through CFG, blending LoRA-s and LoRA-c.
5.2 Comparison with ZipLoRA
Table 6: Comparison with ZipLoRA in two different LoRA setups.
LoRA Setup Methods Composition Quality Image Quality
Character + Style LoRA-S 8.80 8.95
LoRA-C 8.55 9.20
ZipLoRA 9.05 9.40
Character + Object LoRA-S 8.85 9.05
LoRA-C 8.60 9.25
ZipLoRA 8.50 9.10
Although our proposed methods are training-free, we also compare them with the fine-tuning approach ZipLoRA as a baseline for reference. ZipLoRA is based on SDXL and focuses on merging two LoRAs, so we conduct our comparisons under this setup. Specifically, we randomly select publicly available SDXL LoRAs from HuggingFace and create 10 composition sets, combining character + style and character + object for the experiments. All results are compared against the LoRA Merge, and the scores are presented in Table 6.
Since ZipLoRA is specifically designed to merge subject and style LoRAs, it achieves higher scores in the character + style setup compared to our methods, likely due to the benefits of its fine-tuning process. However, in scenarios involving two subjects, such as character + object, our methods demonstrate clear advantages despite being training-free. LoRA-S outperforms in composition quality, while LoRA-C excels in image quality. This indicates that our methods are particularly effective in handling diverse subject combinations, especially when the task involves composing two subjects rather than merging a subject with a style.
6 Conclusion
Broader Impact Statement
Our approaches offer advancements in personalized image generation and customized digital content creation by allowing the combination of arbitrary elements. This capability can be applied to various real-world scenarios, such as virtual try-on and virtual design, leading to positive social impacts. As our method operates in the inference phase and relies solely on the composition of publicly available checkpoints (base models and LoRAs) without requiring additional training, it avoids any negative impact related to model training.
Appendix A Appendix
A.1 Details for Comparative Evaluation using GPT-4V
See Table 7 for the full prompts and Table 8 for a case study of evaluation results.
@@ 3. Object (Huge Two-Handed Burger): two-handed burger, holding a huge burger
A.2 Details for ComposLoRA
See Table 9 for the detailed descriptions of each LoRA in ComposLoRA Testbed.
Table 9: Detailed descriptions of each LoRA in the ComposLoRA.
LoRA Category Trigger Words Source
Anime Style Subset
Zero Suit (Metroid) Clothing zero suit, blue gloves, high heels Link
Hand-drawn Style Style lineart, hand-drawn style Link
Auroral Background Background auroral, starry sky, outdoors Link
Toast Object toast, toast in mouth Link
Realistic Style Subset
Japanese Film Color Style Style film overlay, film grain Link
Bright Style Style bright lighting Link
Forest Background Background slg, river, forest Link
Umbrella Object transparent umbrella Link
Bubble Gum Object blow bubble gum Link